OpenStudy (anonymous):

i need help solving this. please :) y=log(base 5) sqrt x^2-1

OpenStudy (danjs):

solve for what?

OpenStudy (danjs):

you have it solved for y in terms of x already

OpenStudy (anonymous):

i need to find the derivative

OpenStudy (zzr0ck3r):

\(y=\log_5(\sqrt{x^2-1})\)?

OpenStudy (zzr0ck3r):

then we use the chain rule on the log function \(\frac{dy}{dx}=\frac{1}{\ln(5)\sqrt{x^2-1}}*\frac{d}{dx}(\sqrt{x^2-1})\) You with me?

OpenStudy (zzr0ck3r):

@dolphins121

OpenStudy (anonymous):

uhm okay. yeah i think so

OpenStudy (zzr0ck3r):

is this the function \(y=\log_5(\sqrt{x^2-1}) \)

OpenStudy (danjs):

oh

OpenStudy (zzr0ck3r):

that was a question

OpenStudy (zzr0ck3r):

Is this the function \(y=\log_5(\sqrt{x^2-1}) \) ?

OpenStudy (danjs):

just remember \[\log _{b}(x) = \frac{ \ln(x) }{ \ln(b) }\] ln(b) is a constant

OpenStudy (zzr0ck3r):

or even better \((\log_b(x)'=\frac{1}{\ln(b)x}\)

OpenStudy (danjs):

The answer is \[\frac{ x }{ \ln(5)*(x^2 +1) }\] For something to work towards.

OpenStudy (danjs):

\[\frac{ d }{ dx }\log _{5}(\sqrt{x^2 +1}) \]

OpenStudy (zzr0ck3r):

\(\frac{dy}{dx}=\frac{1}{\ln(5)\sqrt{x^2-1}}*\frac{d}{dx}(\sqrt{x^2-1})=\frac{1}{\ln(b)\sqrt{x^2-1}}*\frac{1}{2\sqrt{x^2-1}}*2x=\frac{2x}{2\ln(b)\sqrt{x^2-1}^2}=\\\frac{x}{\ln(b)(x^2-1)} \)

OpenStudy (zzr0ck3r):

does this make sense @dolphins121 ?

OpenStudy (anonymous):

yeah that is the function

OpenStudy (anonymous):

okay so i don't understand where the 2 came from

OpenStudy (zzr0ck3r):

which one? The one on the bottom or the one on the top?

OpenStudy (anonymous):

the one on the bottom

OpenStudy (anonymous):

& when in the second step why did u change ln(5) to ln(b)?

OpenStudy (zzr0ck3r):

the derivative of \(\sqrt{x} \) is \(\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

oh okay

OpenStudy (zzr0ck3r):

\(\sqrt{x}\rightarrow x^\frac{1}{2}\rightarrow \frac{1}{2}x^{\frac{1}{2}-1}=\frac{1}{2x^\frac{1}{2}}\)

OpenStudy (anonymous):

yes i understand now thanks.

OpenStudy (anonymous):

wait uhm so for the final answer why isn't (x^2-1) squared anymore? and why did u take it out of the square root symbol?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

\(\sqrt{x^2-1}*\sqrt{x^2-1}=(\sqrt{x^2-1})^2=x^2-1\)

OpenStudy (anonymous):

ohh okay yeah that makes sense thank u

OpenStudy (zzr0ck3r):

np